OpenStudy (anonymous):

A uniformly charged sphere has a potential on its surface of 460 V. At a radial distance of 32 cm from this surface, the potential is 120 V. What is the radius of the sphere?

OpenStudy (shamim):

do u know the potential on the sphere is \[V=\frac{ 1 }{ 4 \pi \epsilon _{0} }\frac{ q }{ r }\]

OpenStudy (shamim):

q= charge on the sphere

OpenStudy (shamim):

r= radius of sphere

OpenStudy (shamim):

if 32 cm distance frm the sphere , potential is\[V _{1}=\frac{ 1 }{ 4 \pi \epsilon _{0} }\frac{ q }{ r+32 }\]